John likes chicken ribs and chicken wings. Her utility function is U(R,C) =10R2 C. Her weekly 

income is 180 of which she spends exclusively on R and C. The price for the slab of ribs is 20 

shillings and 10 for wings. 

a. What is John’s problem?

b. What is John’s optimal bundle?

c. What is John’s demand function for ribs?

d. Are ribs a normal or inferior good?



1
Expert's answer
2021-03-09T15:19:18-0500

Solution:

a.). What is John’s problem?

Budget constraint:

Price of ribs (R) = 20

Price of wings (C) = 10

180 = 20R + 10C


b.). What is John’s optimal bundle?

U (R, C) = 10R2C

MUR = 20RC

MUC = 10R2

MRSRC = "\frac{MU_{R} }{MU_{C} }" ="\frac{20RC }{10R^{2} }" = "\frac{2C }{R }"


MRS = "\frac{P_{R} }{P_{C} }"


"\frac{2C}{R}" = "\frac{20}{10}"


"\frac{2C}{R}" = 2


C = R

Plug into the budget constraint:

180 = 20R + 10C

180 = 20R + (10"\times"R)

180 = 20R + 10R

180 = 30R

R ="\frac{180}{30}" = 6

R = 6

C = 6

c.). What is John’s demand function for ribs?

180 = 20R + 10C

9 = R + 0.5C

R = 9 – 0.5C

d.). Ribs are a normal good. This is because there is an increased demand for ribs due to rise in consumers demand.